One of your friends wears glasses with a focal length of 36 cm . Assume that the glasses lenses rest 2cm from your friend's eyes.

a) If your friend switches to contact lenses, which rest directly on a person's eyes, what focal length should those lenses have?
Answer #1

your friend is farsighted ( because of +ve focal length )

B)

object distance = 25- 2 = 0.23 m ( 25 cm minimum distance where human eye can focus )

f = 0.36 m

1/v +1/u = 1/f

1/v +1/0.23 = 1/0.36

v≈-0.636923 m

near point = 63.69 + 2 = 65.69 cm

now with contact lens  

object distance u = 0.25 m

image distance = - 0.656923 m

1/- 0.656923 + 1/0.25 = 1/f

from here f≈0.403592 m  

f = 40.359 cm answer

Goodluck for exam Comment in case any doubt, will reply for sure..